Two marbles are drawn on succession from a box contataing 10red, 30white, 20blue & 15 orange marbles with replacement being made after each drawing. Find the probability that
1.both are white
2.first is red & second is white
1
Expert's answer
2012-12-18T08:53:51-0500
1.
Probability of the first white is:30/(10+20+15)=30/45=10/15=2/3
Probability of the second white is: 29/44*2/3=58/132

2.
Probability of the first red is: 10/45=2/9
Probability of the second white is 30/44*2/9=50/44*3=50/132